# Efficient Best-of-Both-Worlds Algorithms for Contextual Combinatorial Semi-Bandits

### v1

@@ 2 Best-of-Both-Worlds Algorithm for Contextual Combinatorial Semi-Bandits
The loss estimation is particularly relevant for the contextual case, as learning the optimal action hinges on a computationally tractable approximation of the unknown parameter $\theta_{t,k}$ governing the loss. Given the covariance matrix $\Sigma_{t,k}=\mathbb{E}[(A_{t})_{k}X_{t}X_{t}^{\top}|\mathcal{F}_{t-1}]$, it is known that we can construct the unbiased estimator $\hat{\theta}_{t,k}$ defined through $\hat{\theta}_{t,k}=\Sigma_{t,k}^{-1}X_{t}\ell_{t}(X_{t},k)(A_{t})_{k}$ for all $k\in[K]$. However, computing this estimator is computationally inefficient as its construction requires computing the inverse of the $d\times d$ covariance matrix $\Sigma_{t,k}$, which is of complexity $\mathcal{O}(d^{3})$. Furthermore, this estimation approach assumes that the covariance matrix is known in advance, which is not the case in most real-world scenarios. To avoid such practical problems, we consider relying on the approach of the Matrix-Geometric-Resampling method proposed by (neu2020efficient, ), as described by the subroutine within Algorithm 1. This approach improves the computational efficiency by order of $\mathcal{O}(d)$ comparing to computing the naïve unbiased estimator and does not require full knowledge of the context distribution $\mathcal{D}$, at the cost of introducing an additional bias in the estimation of the precision matrix $\Sigma_{t,k}^{-1}$. Note that we only require $M_{t}=\lceil 4K\ln(t)/(\alpha_{t}\eta_{t}\lambda_{\min}(\Sigma))\rceil=\widetilde{\mathcal{O}}(\ln(t))$ context samples at every time step. Additionally, an exploration set $E\subseteq\mathcal{A}$ is introduced to bound $\lambda_{\min}(\Sigma_{t,k})$ away from zero. The set $E$ thus must satisfy that, for every $k\in[K]$, there is at least one action $A\in E$ such that $(A)_{k}=1$. For simplicity, we select $E=\big{\{}A\in\{0,1\}^{K}:\sum_{k=1}^{K}(A)_{k}=1\big{\}}$, in which case $|E|=K$. For every $t\in[T]$, we specify the remaining algorithmic parameters as $\eta_{t}=1/\beta_{t}$, where $\beta_{t}=\max\{2,c_{2}\ln T,\beta_{t}^{\prime}\}$ and $\beta_{t+1}^{\prime}=\beta_{t}^{\prime}+c_{1}(1+(m\ln(K/m))^{-1}\sum_{s=1}^{t}H(\bar{A}_{s}(X_{s})))^{-1/2}$. We also let $\alpha_{t}=4K\ln(t)/\lambda_{\min}(\Sigma)$. In addition, we set the problem-dependent constants $c_{1}=\sqrt{(d+\ln T/\lambda_{\min}(\Sigma))K\ln T/(m\ln(K/m))}$ as well as $c_{2}=8K/\lambda_{\min}(\Sigma)$. For initializations, we choose $\beta_{1}^{\prime}=c_{1}\geq 1$. These definitions ensure that $0\leq\alpha_{t}\eta_{t}\leq$ $1/2$ and $0<\eta_{t}\leq 1/2$ throughout $t=1,\ldots,T$ rounds.
@@
    In the adversarial regime, we have
    $\mathcal{R}_{T}=\mathcal{O}\left(\kappa m\sqrt{\ln(K/m)T}\right);$
@@
    In the stochastic regime and the corrupted stochastic regime, we have
    $\mathcal{R}_{T}=\mathcal{O}\left(\frac{\kappa^{2}m^{3/2}\ln((K-m)T)}{\Delta_{\min}}\right).$
@@ 2.1 Regret Analysis
Establishing a best-of-both-worlds guarantee with a computationally efficient algorithm poses several challenges. First, the approach of Zierahn et al. (zierahn2023nonstochastic, ), which uses fixed learning and sampling rates yields only an $\widetilde{\mathcal{O}}(\sqrt{T})$ suboptimal regret bound in the stochastic regime—a result of its constant-learning-rate schedule and the coarse penalty bound in the FTRL analysis. Second, adopting the context-less best-of-both-worlds analysis for combinatorial semi-bandits by Zimmert et al. (zimmert2019beating, ) relies on a hybrid regularizer to control arm-wise entropy; this, however, demands arm-wise bias control that state-of-the-art precision-matrix estimators cannot guarantee under standard assumptions. To overcome these limitations, we exploit the time-varying learning rate to refine the regret analysis to be compatible with the stochastic regime, and we lift the mean-action space (support size $K$) to the space of action distributions (support size exponential in $m$). Together, these approaches enable a self-bounding argument commonly used to achieve a $\widetilde{\mathcal{O}}(\ln T)$ regret bound in the stochastic and corrupted stochastic regimes.
Analyzing regret in the contextual combinatorial semi-bandit setting presents an inherent challenge due to the evolving dependency between the sequence of observed contexts $X_{1},\ldots,X_{T}$, and the learned parameters $\{\tilde{\theta}_{1,k}\}_{k=1}^{K},\ldots,\{\tilde{\theta}_{T,k}\}_{k=1}^{K}$. To address this, we follow a strategy inspired by (neu2020efficient, ; zierahn2023nonstochastic, ), in which we introduce an auxiliary game that simplifies the regret analysis. We define an auxiliary regret notion by introducing a ghost context sample $X_{0}\sim\mathcal{D}$, drawn independently from the data used to construct the estimates $\{\tilde{\theta}_{t,k}\}_{k=1}^{K}$. The regret in this auxiliary game is then
$\widetilde{\mathcal{R}}_{T}(X_{0})=\mathbb{E}\left[\sum_{t=1}^{T}\langle X_{0},\tilde{\theta}_{t,k}\rangle\Big{(}(A_{t})_{k}-\big{(}u^{\star}(X_{0})\big{)}_{k}\Big{)}\right].$
@@
Lemma 2.2 (Original game vs. auxiliary game (neu2020efficient, , Equation (6))).
@@
For all $t\in[T]$, the estimates $\{\tilde{\theta}_{t,k}\}_{k=1}^{K}$ constructed in Algorithm 1 satisfy $\max_{A\in\mathcal{A}}\mathbb{E}\big{[}\sum_{k=1}^{K}\langle x,\hat{\theta}_{t,k}-\tilde{\theta}_{t,k}\rangle(A)_{k}\mid\mathcal{F}_{t-1}\big{]}\leq m/t^{2}$.
@@
$\mathbb{E}\big{[}\widetilde{\mathcal{R}}_{T}(X_{0})\big{]}=L\sqrt{\left(Kd\ln T+\frac{\sqrt{m}K(\ln T)^{2}}{\lambda_{\min}(\Sigma)}\right)\sum_{t=1}^{T}\mathbb{E}[H(\bar{A}_{t}(X_{0}))}]+\frac{8Km\ln(K/m)\ln T}{\lambda_{\min}(\Sigma)},$
@@
Lemma 2.4 serves as the combinatorial bandits analogue of Lemma 3 in (kuroki2024best, ), which was established for linear contextual bandits. However, while (kuroki2024best, ) defines entropy over the space of action distributions, our setting defines entropy over the mean-action polytope. As a result, directly generalizing their entropy bound to settings with multiple arm pulls is insufficient for establishing an optimal regret bound in the stochastic regime for contextual combinatorial bandits. To overcome these limitations, we derive a refined entropy bound by exploiting a careful partitioning of the base arm set $[K]$ and lifting the mean-action space (support size $K$) to the space of action distributions (support size exponential in $m$). This refined bound enables the application of the self-bounding technique from (zimmert2021tsallis, ) in the stochastic setting.
@@
Take a ghost sample $X_{0}\sim\mathcal{D}$ and any action sequence $A_{1},\ldots,A_{t}$ generated under the policy sequence $\pi_{1},\ldots,\pi_{T}$ using Algorithm 1. Suppose that $\sum_{t=1}^{T}\sum_{k:(u^{\star}(X_{0}))_{k}=0}(A_{t})_{k}\geq e$, where we recall that $e$ is the Euler number. The mean-action sequence $\bar{A}_{1},\ldots,\bar{A}_{T}$ generated by Algorithm 1 then satisfies
$\mathbb{E}\left[\sum_{t=1}^{T}H(\bar{A}_{t}(X_{0}))\right]\leq m\ln((K-m)T)\mathbb{E}\left[\sum_{t=1}^{T}\sum_{A\in\mathcal{A}\backslash\{u^{\star}(X_{t})\}}\pi_{t}(A)\right].$
The proof for Theorem 2.1 then follows by combining the insights of all lemmas presented in this section. The complete proof auxiliary results are deferred to Appendix A.1.
@@ 3 Efficient Numerical Scheme for Combinatorial Semi-Bandits
In the adversarial combinatorial semi-bandit setting, as presented in Section 2, the learner must perform FTRL/OSMD updates over the convex hull of the combinatorial action space —an operation that naively entails a $K$-dimensional Bregman projection. By exploiting the KKT conditions of this convex subproblem, we reduce each update to a single one-dimensional root-finding call per round, yielding a more computationally efficient scheme. We first formalize the interaction protocol and then present the OSMD algorithm (Algorithm 2) under the context-free regime. For simplicity, we study in this section the $m$-set setting, where exactly $m$ arms are pulled in each round. Note that in terms of computing Bregman projection, the context-free $m$-set setting is without loss of generality, since Step 4 of Algorithm 1 is computed under a fixed context and we may iterate through $m$ number of $j$-subsets, where $j=1,\ldots,m$. An efficient $\widetilde{\mathcal{O}}(K)$ action-sampling procedure in the $m$-set setting is proposed in (zimmert2019beating, , Appendix B.2).
Interaction Protocol.
Fix the number of base arms $K$ and $m$-set size $m\leq K$. Let $\mathcal{A}=\{A\in\{0,1\}^{K}:\sum_{k=1}^{K}(A)_{k}\leq m\}$. Then, in each round $t=1,\dots,T$, the interaction protocol proceeds as follows. The environment chooses an adversarial loss vector $\ell_{t}=(\ell_{t,1},\dots,\ell_{t,K})\in[-1,1]^{K}$. The learner then plays the vector $A_{t}\in\mathcal{A}$, suffers the total loss $\langle A_{t},y_{t}\rangle=\sum_{k=1}^{K}\ell_{t,k}(A_{t})_{k}$, and observes the coordinate losses $\ell_{t,k}$ for every $k$ such that $(A_{t})_{k}=1$.
Algorithm 2 Online stochastic mirror descent for semi-bandits (lattimore2020bandit, , Algorithm 18)
@@
2:Initialize $\bar{A}_{0}=\arg\min_{a\in\operatorname*{conv}(\mathcal{A})}F(a)$
@@
Given a convex differentiable function $F:\mathcal{A}\to\mathbb{R}$, the Bregman divergence $D_{F}:\operatorname*{conv}(\mathcal{A})\times\operatorname*{conv}(\mathcal{A})\to\mathbb{R}_{+}$ associated with $F$ is defined as
@@
Assuming that the convex potential function $F:\mathcal{A}\to\mathbb{R}$ is separable, that is, $F$ can be expressed as $F(a)=\sum_{k=1}^{K}f(a_{k})$ where $f:\mathbb{R}\to\mathbb{R}$ is convex, and $a_{k}$ is the $k$-th coordinate of $a\in\mathcal{A}$. Hence, the update step amounts to solving the convex subproblem
@@
We derive the KKT conditions as the following. Let us assign the following Lagrangian multipliers $\lambda\in\mathbb{R}$ to the constraint $\sum_{k=1}^{K}a_{k}=m$, then $\mu\in\mathbb{R}^{K}$ to the set of constraints $a_{k}\geq 0,\quad\forall k\in[K]$ and $\nu\in\mathbb{R}^{K}$ to the set of inequality constraints $a_{k}\leq 1,\quad\forall k\in[K]$. Then the Lagrangian has the form $\mathcal{L}(a,\lambda,\mu,\nu)=\sum_{k=1}^{K}[\eta\hat{\ell}_{t,k}a_{k}+f(a_{k})-f((\bar{A}_{t})_{k})-f^{\prime}((\bar{A}_{t})_{k})(a_{k}-(\bar{A}_{t})_{k})]+\lambda(\sum_{k=1}^{K}a_{k}-m)-\sum_{k=1}^{K}\mu_{k}a_{k}+\sum_{k=1}^{K}\nu_{k}(a_{k}-1)$. The resulting KKT conditions for (1) are as follows.
$0\;\leq\;a_{k}\;\leq\;1,\;\textstyle\sum_{k=1}^{K}a_{k}=m,\;\forall k\in[K]$ $(\mathbf{Primal\ feasibility})$
$\mu_{k}\;\geq\;0,\;\nu_{k}\;\geq\;0,\;\forall k\in[K]$ $(\mathbf{Dual\ feasibility})$
$\mu_{k}\,a_{k}=0,\;\nu_{k}\,(a_{k}-1)=0,\;\forall k\in[K]$ $(\mathbf{Complementary\ slackness})$
$\eta\,\hat{\ell}_{t,k}+f^{\prime}(a_{k})-f^{\prime}((\bar{A}_{t})_{k})+\lambda-\mu_{k}+\nu_{k}=0,\;\forall k\in[K]$ $(\mathbf{Stationarity})$
We continue to reformulate the stationarity condition by distinguishing between cases for an arbitrary arm index $k\in[K]$. First, note that when $a_{k}>0$, complementary slackness implies that $\mu_{k}=0$. Substituting into the stationarity condition yields $f^{\prime}(a_{k})+\lambda+c_{k}=0$, where we use the shorthand notation $c_{k}=\eta\hat{\ell}_{t,k}-f^{\prime}((\bar{A}_{t})_{k})$. Inverting $f^{\prime}$ then gives $(f^{\prime})^{-1}(-\lambda-c_{k})=a_{k}$. Next, consider the case when $a_{k}=0$, the stationarity condition together with dual feasibility $\mu_{k}\geq 0$ gives $f^{\prime}(0)+\lambda+c_{k}=\mu_{k}\geq 0$, which implies $(f^{\prime})^{-1}(-\lambda-c_{k})\leq 0$. Since the range of $(f^{\prime})^{-1}$ is $[0,1]$, we conclude that $(f^{\prime})^{-1}(-\lambda-c_{k})=0=a_{k}$. Thus, we have shown that solving the $K$-dimensional convex optimization problem (1) reduces to a one-dimensional root-finding problem
$\sum_{k=1}^{K}(f^{\prime})^{-1}(-\lambda-c_{k})=m.$
Algorithm 3 Bisection algorithm for solving (1)
@@
9:  end if
10:end for
11:return $a_{k}=m/K+(f^{\prime})^{-1}(-\underline{\lambda}-c_{k})-(1/K)\sum_{k=1}^{K}(f^{\prime})^{-1}(-\underline{\lambda}-c_{k})$ for all $k\in[K]$
Similar to the result in (li2024optimism, , Theorem 6.1), which analyzes a perturbation-based algorithm specialized for multi-armed bandits, our bisection algorithm enjoys the following convergence guarantee.
@@
Note that Algorithm 2 calls Algorithm 3 with input $c_{k}=\eta\hat{\ell}_{t,k}-f^{\prime}((\bar{A}_{t})_{k})$ for all $k\in[K]$ in each iteration $t=1,\ldots,T$ in order to compute the mean-action vector $\bar{A}_{t}$. Thus, the width of the search interval $\bar{\lambda}-\underline{\lambda}$ is on the order of $\mathcal{O}(t)$ with high probability. This observation implies that the $t$-th call to Algorithm 3 requires $\mathcal{O}(\ln(\sqrt{K}t/\eta))$ iterations with high probability. In addition, each iteration runs in time $\mathcal{O}(K)$. Hence, if $\eta=\mathcal{O}(\sqrt{T})$, then the $t$-th call of Algorithm 3 runs in time at most $\mathcal{O}(K\ln(\sqrt{KT}))=\widetilde{\mathcal{O}}(K)$ with high probability. Using the sampling scheme of complexity $\widetilde{\mathcal{O}}(K)$ proposed by zimmert2019beating , the efficiency of Algorithm 3 as used by Algorithm 2 is thus comparable to the sampling procedure employed by FTPL (neu2016importance, ).

### v2

@@ 2 Best-of-Both-Worlds Algorithm for Contextual Combinatorial Semi-Bandits
The loss estimation is particularly relevant for the contextual case, as learning the optimal action hinges on a computationally tractable approximation of the unknown parameter $\theta_{t,k}$ governing the loss. Given the covariance matrix $\Sigma_{t,k}=\mathbb{E}[(A_{t})_{k}X_{t}X_{t}^{\top}|\mathcal{F}_{t-1}]$, it is known that we can construct the unbiased estimator $\hat{\theta}_{t,k}$ defined through $\hat{\theta}_{t,k}=\Sigma_{t,k}^{-1}X_{t}\ell_{t}(X_{t},k)(A_{t})_{k}$ for all $k\in[K]$. However, computing this estimator is computationally inefficient as its construction requires computing the inverse of the $d\times d$ covariance matrix $\Sigma_{t,k}$, which is of complexity $\mathcal{O}(d^{3})$. Furthermore, this estimation approach assumes that the covariance matrix is known in advance, which is not the case in most real-world scenarios. To avoid such practical problems, we consider relying on the approach of the Matrix-Geometric-Resampling method proposed by (Neu and Olkhovskaya, 2020), as described by the subroutine within Algorithm 1. This approach improves the computational efficiency by a factor on the order of $\mathcal{O}(d)$ compared to computing the naïve unbiased estimator and does not require full knowledge of the context distribution $\mathcal{D}$, at the cost of introducing an additional bias in the estimation of the precision matrix $\Sigma_{t,k}^{-1}$. Note that we only require $M_{t}=\lceil 4K\ln(t)/(\alpha_{t}\eta_{t}\lambda_{\min}(\Sigma))\rceil=\mathcal{O}(K\ln(T))$ context samples at every time step. Additionally, an exploration set $E\subseteq\mathcal{A}$ is introduced to bound $\lambda_{\min}(\Sigma_{t,k})$ away from zero. The set $E$ thus must satisfy that, for every $k\in[K]$, there is at least one action $A\in E$ such that $(A)_{k}=1$. For simplicity, we select $E=\big\{A\in\{0,1\}^{K}:\sum_{k=1}^{K}(A)_{k}=1\big\}$, in which case $|E|=K$. For every $t\in[T]$, we specify the remaining algorithmic parameters as $\eta_{t}=1/\beta_{t}$, where $\beta_{t}=\max\{2,c_{2}\ln T,\beta_{t}^{\prime}\}$ and $\beta_{t+1}^{\prime}=\beta_{t}^{\prime}+c_{1}(1+(m\ln(K/m))^{-1}\sum_{s=1}^{t}H(\bar{A}_{s}(X_{s})))^{-1/2}$. We also let $\alpha_{t}=4K\ln(t)/\lambda_{\min}(\Sigma)$. In addition, we set the problem-dependent constants $c_{1}=\sqrt{(d+\ln T/\lambda_{\min}(\Sigma))K\ln T/(m\ln(K/m))}$ as well as $c_{2}=8K/\lambda_{\min}(\Sigma)$. For initializations, we choose $\beta_{1}^{\prime}=c_{1}\geq 1$. These definitions ensure that $0\leq\alpha_{t}\eta_{t}\leq$ $1/2$ and $0<\eta_{t}\leq 1/2$ throughout $t=1,\ldots,T$ rounds.
@@
    In the adversarial regime, we have $\mathcal{R}_{T}=\mathcal{O}\left(m\sqrt{K\ln(K/m)T\ln T(d+\ln T/\lambda_{\min}(\Sigma))}\right)$;
@@
    In the stochastic regime and the corrupted stochastic regime, we have $\mathcal{R}_{T}=\mathcal{O}\left(\frac{K\ln Tm^{3/2}\ln((K-m)T)(d+\ln T/\lambda_{\min}(\Sigma))}{\Delta_{\min}}\right)$.
@@ 2.1 Regret Analysis
Establishing a best-of-both-worlds guarantee with a computationally efficient algorithm poses several challenges. First, the approach of Zierahn et al. (Zierahn et al., 2023), which uses fixed learning and sampling rates yields only an $\widetilde{\mathcal{O}}(\sqrt{T})$ suboptimal regret bound in the stochastic regime—a result of its constant-learning-rate schedule and the coarse penalty bound in the FTRL analysis. Second, adopting the context-less best-of-both-worlds analysis for combinatorial semi-bandits by Zimmert et al. (Zimmert et al., 2019) relies on a hybrid regularizer to control arm-wise entropy; this, however, demands arm-wise bias control that state-of-the-art precision-matrix estimators cannot guarantee under standard assumptions. To overcome these limitations, we exploit the time-varying learning rate to refine the regret analysis to be compatible with the stochastic regime, and we lift the mean-action space (support size $K$) to the space of action distributions (support size exponential in $m$). Together, these approaches enable a self-bounding argument commonly used to achieve a $\widetilde{\mathcal{O}}(\ln T)$ regret bound in the stochastic and corrupted stochastic regimes.
Analyzing regret in the contextual combinatorial semi-bandit setting presents an inherent challenge due to the evolving dependency between the sequence of observed contexts $X_{1},\ldots,X_{T}$, and the learned parameters $\{\tilde{\theta}_{1,k}\}_{k=1}^{K},\ldots,\{\tilde{\theta}_{T,k}\}_{k=1}^{K}$. To address this, we follow a strategy inspired by (Neu and Olkhovskaya, 2020; Zierahn et al., 2023), in which we introduce an auxiliary game that simplifies the regret analysis. We define an auxiliary regret notion by introducing a ghost context sample $X_{0}\sim\mathcal{D}$, drawn independently from the data used to construct the estimates $\{\tilde{\theta}_{t,k}\}_{k=1}^{K}$. The regret in this auxiliary game is then
$\widetilde{\mathcal{R}}_{T}(X_{0})=\mathbb{E}\left[\sum_{t=1}^{T}\sum_{k=1}^{K}\langle X_{0},\tilde{\theta}_{t,k}\rangle\Big((A_{t})_{k}-\big(u^{\star}(X_{0})\big)_{k}\Big)\right].$
@@
Lemma 2.2 (Original game vs. auxiliary game (Neu and Olkhovskaya, 2020, Equation (6))).
@@
For all $t\in[T]$, the estimates $\{\tilde{\theta}_{t,k}\}_{k=1}^{K}$ constructed in Algorithm 1 satisfy $\max_{A\in\mathcal{A}}\mathbb{E}\big[\sum_{k=1}^{K}\langle x,\hat{\theta}_{t,k}-\tilde{\theta}_{t,k}\rangle(A)_{k}\mid\mathcal{F}_{t-1}\big]\leq m/t^{2}$ for all $x\in\mathcal{X}$.
@@
$\mathbb{E}\big[\widetilde{\mathcal{R}}_{T}(X_{0})\big]\leq L\sqrt{\left(Kd\ln T+\frac{\sqrt{m}K(\ln T)^{2}}{\lambda_{\min}(\Sigma)}\right)\sum_{t=1}^{T}\mathbb{E}[H(\bar{A}_{t}(X_{0}))}]+\frac{8Km\ln(K/m)\ln T}{\lambda_{\min}(\Sigma)},$
@@
Lemma 2.4 serves as the combinatorial bandits analogue of Lemma 3 in (Kuroki et al., 2024), which was established for linear contextual bandits. However, while (Kuroki et al., 2024) defines entropy over the space of action distributions, our setting defines entropy over the mean-action polytope. As a result, directly generalizing their entropy bound to settings with multiple arm pulls is insufficient for establishing an optimal regret bound in the stochastic regime for contextual combinatorial bandits. To overcome these limitations, we derive a refined entropy bound by exploiting a careful partitioning of the base arm set $[K]$ and lifting the mean-action space (support size $K$) to the space of action distributions (support size exponential in $m$). This refined bound enables the application of the self-bounding technique from (Zimmert and Seldin, 2021) in the stochastic setting.
@@
Take a ghost sample $X_{0}\sim\mathcal{D}$ and any action sequence $A_{1},\ldots,A_{T}$ generated under the policy sequence $\pi_{1},\ldots,\pi_{T}$ using Algorithm 1. Suppose that $\sum_{t=1}^{T}\sum_{k:(u^{\star}(X_{0}))_{k}=0}(A_{t})_{k}\geq e$, where $e$ denotes Euler’s number. The mean-action sequence $\bar{A}_{1},\ldots,\bar{A}_{T}$ generated by Algorithm 1 then satisfies
$\mathbb{E}\left[\sum_{t=1}^{T}H(\bar{A}_{t}(X_{0}))\right]\leq m\ln((K-m)T)\mathbb{E}\left[\sum_{t=1}^{T}\sum_{A\in\mathcal{A}\backslash\{u^{\star}(X_{t})\}}\pi_{t}(A|X_{t})\right].$
The proof for Theorem 2.1 then follows by combining the insights of all lemmas presented in this section. The complete proof and auxiliary results are deferred to Appendix A.1.
@@ 3 Efficient Numerical Scheme for Combinatorial Semi-Bandits
In the adversarial combinatorial semi-bandit setting, as presented in Section 2, the learner must perform FTRL/OSMD updates over the convex hull of the combinatorial action space—an operation that naively entails a $K$-dimensional Bregman projection. By exploiting the KKT conditions of this convex subproblem, we reduce each update to a single one-dimensional root-finding call per round, yielding a more computationally efficient scheme. We first formalize the interaction protocol and then present the OSMD algorithm (Algorithm 2) under the context-free regime. For simplicity, we study in this section the $m$-set setting, where exactly $m$ arms are pulled in each round. Note that in terms of computing Bregman projection, the context-free $m$-set setting is without loss of generality, since Step 4 of Algorithm 1 is computed under a fixed context and we may iterate through $m$ number of $j$-subsets, where $j=1,\ldots,m$. An efficient $\widetilde{\mathcal{O}}(K)$ action-sampling procedure in the $m$-set setting is proposed in (Zimmert et al., 2019, Appendix B.2).
Interaction Protocol. Fix the number of base arms $K$ and $m$-set size $m\leq K$. Let $\mathcal{A}=\{A\in\{0,1\}^{K}:\sum_{k=1}^{K}(A)_{k}{=}m\}$. Then, in each round $t=1,\dots,T$, the interaction protocol proceeds as follows. The environment chooses an adversarial loss vector $\ell_{t}=(\ell_{t,1},\dots,\ell_{t,K})\in[-1,1]^{K}$. The learner then plays the vector $A_{t}\in\mathcal{A}$, suffers the total loss $\langle A_{t},\ell_{t}\rangle=\sum_{k=1}^{K}\ell_{t,k}(A_{t})_{k}$, and observes the coordinate losses $\ell_{t,k}$ for every $k$ such that $(A_{t})_{k}=1$.
Algorithm 2 Online stochastic mirror descent for semi-bandits (Lattimore and Szepesvári, 2020, Algorithm 18)
@@
2:Initialize $\bar{A}_{1}=\arg\min_{a\in\operatorname*{conv}(\mathcal{A})}F(a)$
@@
Given a convex differentiable function $F:\operatorname*{conv}(\mathcal{A})\to\mathbb{R}$, the Bregman divergence $D_{F}:\operatorname*{conv}(\mathcal{A})\times\operatorname*{conv}(\mathcal{A})\to\mathbb{R}_{+}$ associated with $F$ is defined as
@@
Assuming that the convex potential function $F:\operatorname*{conv}(\mathcal{A})\to\mathbb{R}$ is separable, that is, $F$ can be expressed as $F(a)=\sum_{k=1}^{K}f(a_{k})$ where $f:\mathbb{R}\to\mathbb{R}$ is convex, and $a_{k}$ is the $k$-th coordinate of $a\in\mathcal{A}$. Hence, the update step amounts to solving the convex subproblem
@@
We derive the KKT conditions as the following. Let us assign the following Lagrangian multipliers $\lambda\in\mathbb{R}$ to the constraint $\sum_{k=1}^{K}a_{k}=m$, then $\mu\in\mathbb{R}^{K}$ to the set of constraints $a_{k}\geq 0,\quad\forall k\in[K]$ and $\nu\in\mathbb{R}^{K}$ to the set of inequality constraints $a_{k}\leq 1,\quad\forall k\in[K]$. Then the Lagrangian has the form $\mathcal{L}(a,\lambda,\mu,\nu)=\sum_{k=1}^{K}[\eta\hat{\ell}_{t,k}a_{k}+f(a_{k})-f((\bar{A}_{t})_{k})-f^{\prime}((\bar{A}_{t})_{k})(a_{k}-(\bar{A}_{t})_{k})]+\lambda(\sum_{k=1}^{K}a_{k}-m)-\sum_{k=1}^{K}\mu_{k}a_{k}+\sum_{k=1}^{K}\nu_{k}(a_{k}-1)$. The resulting KKT conditions for equation 1 are as follows.
$0\;\leq\;a_{k}\;\leq\;1,\;\textstyle\sum_{k=1}^{K}a_{k}=m,\;\forall k\in[K]$ $(\text{Primal\ feasibility})$
$\mu_{k}\;\geq\;0,\;\nu_{k}\;\geq\;0,\;\forall k\in[K]$ $(\text{Dual\ feasibility})$
$\mu_{k}\,a_{k}=0,\;\nu_{k}\,(a_{k}-1)=0,\;\forall k\in[K]$ $(\text{Complementary\ slackness})$
$\eta\,\hat{\ell}_{t,k}+f^{\prime}(a_{k})-f^{\prime}((\bar{A}_{t})_{k})+\lambda-\mu_{k}+\nu_{k}=0,\;\forall k\in[K]$ $(\text{Stationarity})$
We continue to reformulate the stationarity condition by distinguishing between cases for an arbitrary arm index $k\in[K]$. First, note that when $a_{k}>0$, complementary slackness implies that $\mu_{k}=0$. Substituting into the stationarity condition yields $f^{\prime}(a_{k})+\lambda+c_{k}=0$, where we use the shorthand notation $c_{k}=\eta\hat{\ell}_{t,k}-f^{\prime}((\bar{A}_{t})_{k})$. Inverting $f^{\prime}$ then gives $(f^{\prime})^{-1}(-\lambda-c_{k})=a_{k}$. Next, consider the case when $a_{k}=0$, the stationarity condition together with dual feasibility $\mu_{k}\geq 0$ gives $f^{\prime}(0)+\lambda+c_{k}=\mu_{k}\geq 0$, which implies $(f^{\prime})^{-1}(-\lambda-c_{k})\leq 0$. Since the range of $(f^{\prime})^{-1}$ is $[0,1]$, we conclude that $(f^{\prime})^{-1}(-\lambda-c_{k})=0=a_{k}$.
Thus, we have shown that solving the $K$-dimensional convex optimization problem equation 1 reduces to a one-dimensional root-finding problem $\sum_{k=1}^{K}(f^{\prime})^{-1}(-\lambda-c_{k})=m.$
Algorithm 3 Bisection algorithm for solving equation 1
@@
9:return $a_{k}=m/K+(f^{\prime})^{-1}(-\underline{\lambda}-c_{k})-(1/K)\sum_{j=1}^{K}(f^{\prime})^{-1}(-\underline{\lambda}-c_{j})$ for all $k\in[K]$
Similar to the result in (Li et al., 2024, Theorem 6.1), which analyzes a perturbation-based algorithm specialized for multi-armed bandits, our bisection algorithm enjoys the following convergence guarantee.
@@
Note that Algorithm 2 calls Algorithm 3 with input $c_{k}=\eta\hat{\ell}_{t,k}-f^{\prime}((\bar{A}_{t})_{k})$ for all $k\in[K]$ in each iteration $t=1,\ldots,T$ in order to compute the mean-action vector $\bar{A}_{t}$. Thus, the width of the search interval $\bar{\lambda}-\underline{\lambda}$ is on the order of $\mathcal{O}(t)$ with high probability. This observation implies that the $t$-th call to Algorithm 3 requires $\mathcal{O}(\ln(\sqrt{K}\eta T))$ iterations with high probability. In addition, each iteration runs in time $\mathcal{O}(K)$. Hence, if $\eta=\mathcal{O}(1/\sqrt{T})$, then the $t$-th call of Algorithm 3 runs in time at most $\mathcal{O}(K\ln(\sqrt{KT}))=\widetilde{\mathcal{O}}(K)$ with high probability. Using the sampling scheme of complexity $\widetilde{\mathcal{O}}(K)$ proposed by Zimmert et al. (2019), the efficiency of Algorithm 3 as used by Algorithm 2 is thus comparable to the sampling procedure employed by FTPL (Neu and Bartók, 2016).
We remark on the extension of our results in this section to settings beyond $m$-set such as a partition matroid (Oxley, 2011). Although we now have cardinality constraints per partition instead of a single cardinality constraint, each partition is handled by its own scalar root-finding problem, and every arm enters exactly one partition. Hence, the work across all partitions sums to $K$, which is the same as the uniform-matroid case. Building on the previous observation, the Bregman projection using our method still runs in time $O(K\ln(1/\varepsilon))$; the bisection method over partition $i$ runs in time $O(c_{i}\ln(1/\varepsilon))$ with $K=\sum_{i}c_{i}$.